# RelayAttention for Efficient Large Language Model Serving with Long System Prompts

### v2

@@ Abstract
Practical large language model (LLM) services may involve a long system prompt, which specifies the instructions, examples, and knowledge documents of the task and is reused across numerous requests. However, the long system prompt causes throughput/latency bottlenecks as the cost of generating the next token grows w.r.t. the sequence length. This paper aims to improve the efficiency of LLM services that involve long system prompts. Our key observation is that handling these system prompts requires heavily redundant memory accesses in existing causal attention computation algorithms. Specifically, for batched requests, the cached hidden states (i.e., key-value pairs) of system prompts are transferred from off-chip DRAM to on-chip SRAM multiple times, each corresponding to an individual request. To eliminate such a redundancy, we propose RelayAttention, an attention algorithm that allows reading these hidden states from DRAM exactly once for a batch of input tokens. RelayAttention is a free lunch: it maintains the generation quality while requiring no model retraining, as it is based on a mathematical reformulation of causal attention. Code is available at https://github.com/rayleizhu/vllm-ra.
@@
[…]
${}^{\dagger}$ Corresponding author.
@@ 1 Introduction
LLM services commonly use an application-specific system prompt OpenAI (2023a) to specify the task’s instructions. The system prompt is concatenated with the user prompt as the full input to the LLM for response generation and is shared by all requests to a service. The system prompt becomes long if the service provider wants to provide detailed guidelines and examples for better response quality or apply more restrictions/policies for ethical safety. As the sequence length that LLMs can process grows Anthropic (2023); Chen et al. (2023b); DeepSeek-AI et al. (2024), some emerging professional applications, such as legal analysis Cui et al. (2023); Nay et al. (2023), healthcare applications Steinberg et al. (2021); Rasmy et al. (2021), and the shopping assistant example shown in Fig. 2, may include one or more knowledge documents to provide domain-specific knowledge, resulting in even longer system prompts. Although long system prompts are beneficial to improving the generation quality or enabling new applications, they also pose a challenge to the LLM service: the inference throughput and latency of the service can be heavily degraded, thus increasing the per-request cost. This is inherently caused by the causal attention, in which each new token is generated by “looking at” all precedent ones.
@@
In this paper, we propose a novel approach to mitigate the efficiency problem of using long system prompts in LLM services. Our key observation is that there are not only redundant memory footprint Kwon et al. (2023) and computations Gim et al. (2023) corresponding to the system prompt, but also unnecessary memory accesses during causal attention computation. Specifically, while the system prompt is shared by all requests, its hidden states (i.e., key-value pairs) are read from DRAM multiple times by existing attention algorithms such as PagedAttention Kwon et al. (2023) and FlashAttention Dao et al. (2022); Dao (2023), each for an individual request in the batch. This severely slows down LLM inferences, which are known to be memory-bound (Section 3.2). To eliminate such redundant memory access, we propose RelayAttention, an exact algorithm to compute causal attention based on a mathematical reformulation of it. The key idea of RelayAttention is to group the matrix-vector multiplications corresponding to the system prompt into matrix-matrix multiplications, which allow loading the hidden states of the system prompt from DRAM exactly once for all request tokens in a batch (Section 3.3). We provide an in-depth analysis of the theoretic speedup via redundancy reduction with IO-awareness (Section 3.4). Our empirical results further verify the efficiency: integrating RelayAttention into vLLM Kwon et al. (2023), an already highly optimized production-level LLM serving system, we still observe up to $\times2$ sustainable request rate and 2.0×2.0$\times$2.0 × throughput with the Llama2-7B model for a chatbot workload. Similar efficiency improvements are also observed for several other popular LLMs and are consistent on several data center GPUs. The efficiency gains continue growing with longer system prompts.
@@ 3.3 LLM Serving with RelayAttention
This rescaling can also be applied to the second term on the right side of Eq. 5, and thus $\mathbf{o}_{t}$ is a convex combination of two scaled attention terms:
@@
Peripheral adaptations. There are two major adaptations needed to make RelayAttention work better within existing inference systems. First, instead of using a single KV cache for both the system prompt and the request-specific context, we use a separate system KV cache to store system KVs and fill it offline before model serving. This can be viewed as a combination of prefix sharing in PagedAttention, which eliminates redundant memory footprint of system KVs, and PromptCache Gim et al. (2023), which eliminates redundant computation in the prompt phase. Second, as the system KVs are already computed offline, we add an offset of $s$ (i.e., the length of the system prompt) in the position of those request-specific context tokens to make sure of correct position embedding.
@@ 3.4 Theoretical Speedup
$n^{\prime}=\underbrace{(bd+sd+bd)}_{\text{system attetion}}+\underbrace{(bd+bcd+bd)}_{\text{context attention}}+\underbrace{3bd}_{\text{relay fusion}}.$ (9)
@@
In Fig. 5, we plot the speedup brought by using RelayAttention. The small gap between the practical and theoretical curves verifies our analysis.
Figure 5: The theoretical and practical speedups for casual attention computation with RelayAttention. We set the batch size to 32 and context length to 128, and plot the speedup w.r.t. the length of the system prompt. A40 GPU is used.
@@ 4.1 Experimental Setup
  • vLLM  [^2] https://github.com/vllm-project/vllm: a state-of-the-art open-source LLM inference system designed for high throughput LLM serving. Note that the core component of vLLM, PagedAttention Kwon et al. (2023), allows storing the shared system KVs exactly once by the prefix sharing technique mentioned in their paper, but this technique is not included in the public code release. Considering the importance to save memory for a higher concurrency, we implement a stronger baseline, vLLM-PS as specified below.
@@ 4.2 Noninteractive Batch Processing
We plot the throughputs w.r.t. the length of system prompt for processing ShareGPTv3 on the three GPUs in Fig. 6. For vLLM, the throughputs degrade heavily as the system prompt becomes long for two reasons: (1) the system prompt occupies too much memory, and thus heavily limits the batch size/concurrency of decoding; (2) it takes too much time to handle long system prompts during causal attention computation. With prefixing sharing, vLLM-PS solves the first problem and achieves up to 108% improvement on the throughput. Our vLLM-RL further solves the second problem and increases the throughput to 1.06×1.06$\times$1.06 × to 4.36×4.36$\times$4.36 × of vLLM. Table 3 shows results of the few-shot test on MMLU. We can see that using a long system prompt to include more examples is crucial for improving accuracy. In the case of the 5-shot test, our vLLM-RA provides a 76% reduction of processing time on both A40 and A100-SXM4-80GB GPUs.
@@ 4.4 The Improvement for More Models
5 Limitations and Future Work
The limitations of RelayAttention can be reflected by the theoretical speedup (Eq. 10). First, it helps batched inference ($b>1$). The larger the batch size, the more efficient RelayAttention is. When there is only one request, which is the typical case on device-side applications, RelayAttention does not help. Therefore, RelayAttention is suitable for cloud-serving scenarios. Second, when the request-specific contexts (including user prompts and responses) are long (e.g., 2×2$\times$2 × longer than the shared system prompt), the computation time is dominated by the processing of them; thus the efficiency gain will diminish. However, as the context length has a long-tailed distribution in many applications (e.g., chatbots), where the majority of user prompts and responses are short, the efficiency gain brought by RelayAttention is still considerable. In future work, we will explore more applications by customizing a LLM with long system prompts.
@@ Appendix A Implementation of RelayAttention
Appendix B System Level Design of vLLM-RA
As mentioned in Section 3.3, it is easy to integrate RelayAttention into existing inference system with the replacement of attention computation function and several peripheral adaptations. In Fig. 9, we summarize the system level design of vLLM-RA in a comparison with vLLM.
Figure 9: Comparison of the system level design of vLLM (top) and vLLM-RA (bottom). The modifications of vLLM-RA are highlighted in red.
Appendix C More Information of The Datasets
@@
Appendix D Benchmark with Synthetic Workloads

### v3

@@ Abstract
A practical large language model (LLM) service may involve a long system prompt, which specifies the instructions, examples, and knowledge documents of the task and is reused across requests. However, the long system prompt causes throughput/latency bottlenecks as the cost of generating the next token grows w.r.t. the sequence length. This paper aims to improve the efficiency of LLM services that involve long system prompts. Our key observation is that handling these system prompts requires heavily redundant memory accesses in existing causal attention computation algorithms. Specifically, for batched requests, the cached hidden states (i.e., key-value pairs) of system prompts are transferred from off-chip DRAM to on-chip SRAM multiple times, each corresponding to an individual request. To eliminate such a redundancy, we propose RelayAttention, an attention algorithm that allows reading these hidden states from DRAM exactly once for a batch of input tokens. RelayAttention is a free lunch: it maintains the generation quality while requiring no model retraining, as it is based on a mathematical reformulation of causal attention. We have observed significant performance improvements to a production-level system, vLLM, through integration with RelayAttention. The improvements are even more profound with longer system prompts.
@@
Lei Zhu1  Xinjiang Wang2  Wayne Zhang2  Rynson Lau1† 1 City University of Hong Kong   2 SenseTime Research lzhu68-c@my.cityu.edu.hk, {wangxinjiang,wayne.zhang}@sensetime.com, Rynson.Lau@cityu.edu.hk Code: https://github.com/rayleizhu/vllm-ra
Corresponding author.
@@ 1 Introduction
LLM services commonly use an application-specific system prompt OpenAI (2023a) to specify the task’s instructions. The system prompt is concatenated with the user prompt as the full input to the LLM for response generation and is shared by all requests to a service. The system prompt becomes long if the service provider wants to provide detailed guidelines and examples for better response quality or apply more restrictions/policies for ethical safety. As the sequence length that LLMs can process grows Anthropic (2023); Chen et al. (2023b); Bi et al. (2024), some emerging professional applications, such as legal analysis Cui et al. (2023); Nay et al. (2023), healthcare applications Steinberg et al. (2021); Rasmy et al. (2021), and the shopping assistant example shown in Fig. 2, may include one or more knowledge documents to provide domain-specific knowledge, resulting in even longer system prompts. Although long system prompts are beneficial to improving the generation quality or enabling new applications, they also pose a challenge to the LLM service: the inference throughput and latency of the service can be heavily degraded, thus increasing the per-request cost. This is inherently caused by the causal attention, in which each new token is generated by “looking at” all precedent ones.
@@
In this paper, we propose a novel approach to mitigate the efficiency problem of using long system prompts in LLM services. Our key observation is that there are not only redundant memory footprint Kwon et al. (2023) and computations Gim et al. (2023) corresponding to the system prompt, but also unnecessary memory accesses during causal attention computation. Specifically, while the system prompt is shared by all requests, its hidden states (i.e., key-value pairs) are read from DRAM multiple times by existing attention algorithms such as PagedAttention Kwon et al. (2023) and FlashAttention Dao et al. (2022); Dao (2023), each for an individual request in the batch. This severely slows down LLM inferences, which are known to be memory-bound (Section 3.2). To eliminate such redundant memory access, we propose RelayAttention, an exact algorithm to compute causal attention based on a mathematical reformulation of it. The key idea of RelayAttention is to group the matrix-vector multiplications corresponding to the system prompt into matrix-matrix multiplications, which allow loading the hidden states of the system prompt from DRAM exactly once for all request tokens in a batch (Section 3.3). As a result, the attention inference latency grows much slower than PagedAttention w.r.t. the length of system prompt, as shown in Fig. 1. We provide an in-depth analysis of the theoretic speedup of the standalone attention based on the IO redundancy reduction (Section 3.4). Our empirical results for end-to-end serving further verify the efficiency: integrating RelayAttention into vLLM Kwon et al. (2023), an already highly optimized production-level LLM serving system, we still observe up to $\times2$ sustainable request rate and 2.0×2.0$\times$2.0 × throughput with the Llama2-7B model for a chatbot workload. Similar efficiency improvements are also observed for several other popular LLMs and are consistent on several data center GPUs. The efficiency gains continue growing with longer system prompts.
@@ 3.3 LLM Serving with RelayAttention
This rescaling trick Milakov and Gimelshein (2018); Rabe and Staats (2021) can also be applied to the second term on the right side of Eq. 5, and thus $\mathbf{o}_{t}$ is a convex combination of two scaled attention terms:
@@
Figure 5: Key modifications (high-lighted in red in the bottom) to integrate RelayAttention into an existing LLM serving system (top).
Peripheral adaptations. There are two major adaptations needed to make RelayAttention work better within existing inference systems. First, instead of using a single KV cache for both the system prompt and the request-specific context, we use a separate system KV cache to store system KVs and fill it offline before model serving. This can be viewed as a combination of prefix sharing in PagedAttention, which eliminates redundant memory footprint of system KVs, and PromptCache Gim et al. (2023), which eliminates redundant computation in the prompt phase. Second, as the system KVs are already computed offline, we add an offset of $s$ (i.e., the length of the system prompt) in the position of those request-specific context tokens to make sure of correct position embedding. Fig. 5 summarizes the key modifications to integrate RelayAttention into an existing LLM serving system (e.g., vLLM Kwon et al. (2023) or TensorRT-LLM Nvidia (2023)).
@@ 3.4 Theoretical Speedup
Figure 6: The theoretical and practical speedups for a standalone Llama-30B casual attention with RelayAttention. We plot the speedup w.r.t. the length of the system prompt under two context lengths (128 and 256) and four batch sizes (4, 8, 16, and 32), on an A40 GPU.
@@
$n^{\prime}=\underbrace{(bd+sd+bd)}_{\text{system attention}}+\underbrace{(bd+bcd+bd)}_{\text{context attention}}+\underbrace{3bd}_{\text{relay fusion}}.$ (9)
@@
In Fig. 6, we plot the speedup brought by using RelayAttention. The small gaps between the practical and corresponding theoretical curves verify our analysis.
@@ 4.1 Experimental Setup
  • vLLM  [^2] https://github.com/vllm-project/vllm: a state-of-the-art open-source LLM inference system designed for high throughput LLM serving. We use the v0.2.6 release. Note that the core component of vLLM, PagedAttention Kwon et al. (2023), allows storing the shared system KVs exactly once by the prefix sharing technique mentioned in their paper, but this technique is not included in the vLLM v0.2.6 release. Considering the importance to save memory for a higher concurrency, we implement a stronger baseline, vLLM-PS as specified below.
@@
Note that, vLLM and the two variants mentioned above use dynamic batch size: requests are dynamically batched with varied batch sizes in {1, 2, 4, 8, 16, 24, 32, 40, …, 256} by a scheduler, depending on the available memory and the number of pending requests in the queue. vLLM would schedule as many requests as possible each time to maximize the hardware utilization.
@@ 4.2 Noninteractive Batch Processing
We plot the throughputs w.r.t. the length of system prompt for processing ShareGPTv3 on the three GPUs in Fig. 7. For vLLM, the throughputs degrade heavily as the system prompt becomes long for two reasons: (1) the system prompt occupies too much memory, and thus heavily limits the batch size/concurrency of decoding; (2) it takes too much time to handle long system prompts during causal attention computation. With prefixing sharing, vLLM-PS solves the first problem and achieves up to 108% improvement on the throughput. Our vLLM-RL further solves the second problem and increases the throughput from 1.06×1.06$\times$1.06 × to 4.36×4.36$\times$4.36 × of vLLM. Table 3 shows results of the few-shot test on MMLU. We can see that using a long system prompt to include more examples is crucial for improving accuracy. In the case of the 5-shot test, our vLLM-RA provides a 76% reduction of processing time on both A40 and A100-SXM4-80GB GPUs.
@@ 5 Conclusion
Limitations
The limitations of RelayAttention can be reflected by the theoretical speedup (Eq. 10). First, it helps batched inference ($b>1$). The larger the batch size, the more efficient RelayAttention is. When there is only one request, which is the typical case on device-side applications, RelayAttention does not help. Therefore, RelayAttention is suitable for cloud-serving scenarios. Second, when the request-specific contexts (including user prompts and responses) are long (e.g., 2×2$\times$2 × longer than the shared system prompt), the computation time is dominated by the processing of them; thus the efficiency gain will diminish. However, as the context length has a long-tailed distribution in many applications (e.g., chatbots), where the majority of user prompts and responses are short, the efficiency gain brought by RelayAttention is still considerable.
Broader Impact Statement
It is revealed by RelayAttention that the inference cost caused by a shared prefix can be amortized by a batch of requests. As a result, the shared prefix is much cheaper than request-specific contexts. This can reduce the hosting cost of existing LLM applications. It may also encourage LLM customization with longer system prompts or prefix-tuning Li and Liang (2021) in practice.
@@ Appendix A Implementation of RelayAttention
Appendix B More Information of The Datasets
@@
Appendix C Benchmark with Synthetic Workloads
@@
Appendix D Extension to Multi-Application Hosting
RelayAttention assumes that incoming requests share the same system prompt, which implies the serving process provides only one application. Although this is reasonable for applications (e.g., ChatGPT) with wide usage, deploying multiple applications in a single serving process is more economical in some scenarios. In such cases, it just requires some engineering efforts to support: for example, tagging each request with an application ID and implementing a scheduler to batch requests with the same application ID each time.